OpenStudy (anonymous):

can somebody plzs help f(x)=3x-9 and g(x)=x^2 wht is (g*f)(5)

OpenStudy (hanner_b_nanner):

plug in 5 for x... :))

OpenStudy (anonymous):

gof =(3x-9)^2 now for 5 gof(5)=(3*5-9)^2 gof(5)=(15-9)^2 gof(5)=(6)^2 gof(5)=36